# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 6. The Other Direction: From Standard Isoperimetry to Log-Sobolev
We now prove the other direction (i.e., the upper bound) in Theorem 1. Recall $\xi$ defined in (5.8) and $\bar{\Lambda}$ defined around (1.5).
@@
Under Assumption 1, for any $\alpha>0$,
(6.1) $\bar{\Lambda}(\alpha)$ $\leq\lim_{t\downarrow\alpha}\xi(t)\leq\sqrt{\breve{\Theta}(\alpha)}.$
@@
We first prove the first inequality in (6.1). Let $A$ be a set approximately attaining $\Gamma_{n}(e^{-n\alpha},n\tau)$ for $\nu=\nu_{X}=\nu_{Y}$ (given in (3.2)) within a factor $1-\epsilon$ where $\epsilon>0$, i.e., $A$ has probability $e^{-n\alpha}$ and $\nu_{n}(A^{\sqrt{n\tau}})\geq(1-\epsilon)\Gamma_{n}(e^{-n\alpha},n\tau)$. Let $r:=\sqrt{n\tau}$. Since $f(t):=\ln\nu_{n}(A^{t})$ is monotone, it is differentiable almost everywhere on $(0,r)$. As a basic property of monotone functions, $f^{\prime}$ is integrable over $[0,r]$ and $\int_{0}^{r}f^{\prime}(t){\text{\rm d}}t\leq f(r)-f(0)$ (see e.g., [60, Corollary 4 on p.113]), which implies that there is some $r_{0}\in(0,r)$ such that $\frac{f(r)-f(0)}{r}\geq f^{\prime}(r_{0}),$i.e.,
$\frac{\ln[\nu_{n}(A^{r})/\nu_{n}(A)]}{r}\geq\frac{\nu_{n}^{+}(A^{r_{0}})}{\nu_{n}(A^{r_{0}})}.$
@@
$\frac{\ln[\nu_{n}(A^{r})/\nu_{n}(A)]}{r}\geq\frac{I_{n}(\nu_{n}(A^{r_{0}}))}{\nu_{n}(A^{r_{0}})}\geq\frac{I_{n}(\nu_{n}(A^{r}))}{\nu_{n}(A^{r})}.$
@@
$\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$ $\geq\lim_{n\to\infty}\frac{\ln[\nu_{n}(A^{\sqrt{n\tau}})/\nu_{n}(A)]}{n\sqrt{\tau}}\geq\limsup_{n\to\infty}\frac{I_{n}(a_{n,\tau})}{a_{n,\tau}\sqrt{n}},$
where $a_{n,\tau}=\nu_{n}(A^{\sqrt{n\tau}})$.
Since $A$ approximately attains $\Gamma_{n}(e^{-n\alpha},n\tau)$ within a factor $1-\epsilon$, by Statement 2 in Theorem 8 again, it holds that $\liminf_{n\to\infty}-\frac{1}{n}\ln a_{n,\tau}\geq\psi(\alpha,\tau).$ Hence, for any $\delta>0$, $\bar{\Lambda}(\psi(\alpha,\tau)-\delta)\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$
We can assume that $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$, since otherwise, $\xi(\alpha)=\infty$ and nothing is needed to prove. Hence, given any $\delta>0$, it holds that $\psi(\alpha,\tau)>\alpha-\delta.$ for all sufficiently small $\tau$, which, combined with the inequality above, implies $\bar{\Lambda}(\alpha-\delta)\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$ Letting $\tau\downarrow 0$ yields that
(6.2) $\bar{\Lambda}(\alpha-\delta)$ $\leq\xi(\alpha).$
Since $\delta>0$ is arbitrary, we obtain the first inequality in (6.1).
We next prove the second inequality in (6.1). Substituting the expression of $\psi(\alpha,\tau)$ into the bound above yields that
$\xi(\alpha)$ $=\liminf_{\tau\downarrow 0}\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$
@@
Since $\pi_{Y|W=i}=\pi_{X|W=i}$ is a feasible solution to the supremum above, we have $\beta_{i}\leq\alpha_{i},i\in\{0,1\}$.
@@
We now need the HWI inequality for the Riemannian manifold setting [65, Corollary 20.13]. The original version for Euclidean spaces is due to Otto and Villani [54].
Theorem 17 (HWI Inequality [65, Corollary 20.13]).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that Assumption 1 holds. Then, for any two probability measures $\mu_{0}=\rho_{0}\nu,\mu_{1}=\rho_{1}\nu$ in $P_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\mu_{1}\|\nu)<\infty$ and $\rho_{0}$ is Lipschitz, it holds that
@@
This theorem implies $|\partial^{-}D|(\mu\|\nu)\leq\sqrt{I(\mu\|\nu)}.$ In fact, this is an equality; see [6, Theorem 9.3]. We restrict $\pi_{X|W=i},i\in\{0,1\}$ to be in $P_{2}^{\mathrm{ac}}(\mathcal{M})$. Therefore,
$\xi(\alpha)$ $\leq\inf_{\begin{subarray}{c}\pi_{XW}:\pi_{X|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),\forall i,\\ D(\pi_{X|W}\|\nu|\pi_{W})=\alpha\end{subarray}}\sqrt{I(\pi_{X|W}\|\nu|\pi_{W})}=\sqrt{\breve{\Theta}(\alpha)}.$
Combining this with (6.2) yields that $\bar{\Lambda}(\alpha-\delta)\leq\xi(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$ That is, for $\alpha,\delta>0$, $\bar{\Lambda}(\alpha)\leq\xi(\alpha+\delta)\leq\sqrt{\breve{\Theta}(\alpha+\delta)}.$ Since $\breve{\Theta}(t)$ is continuous in $t>0$, letting $\delta\downarrow 0$ yields $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$
@@
By an approximation argument, we may assume that the probability measures $\mu$ in the nonlinear log-Sobolev inequality have smooth density $f:={\text{\rm d}}\mu/{\text{\rm d}}\nu$. Obviously, $\int f\,d\nu=1$. Let $f^{\otimes n}$ be the $n$-fold product of $f$ with itself, which satisfies ${\text{\rm d}}\mu_{n}/{\text{\rm d}}\nu_{n}=f^{\otimes n}$ with $\mu_{n}=\mu^{\otimes n}$. Then, the coarea formula (see e.g., [19, Exercise III.12]) states:
@@
$\Theta_{k}(\alpha):=\inf_{\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}):D(\mu\|\nu^{\otimes k})\geq\alpha}I(\mu\|\nu^{\otimes k}).$
@@
Appendix A: Proof of Theorem 2
@@
Appendix A-1. Inequality (b) in Statement 1
We first prove Inequality (b) in Statement 1 of Theorem 2. Due to Theorem 12, or more precisely, due to (3.1) and (4.2) in [45] but with $K_{CD}\geq 0$, it holds that for any $n$,
@@ Appendix A-1. Inequality (b) in Statement 1
Appendix A-2. Inequality (l)
Inequality (l) is exactly the following inequality: for any $n\geq 1$,

### v4

@@ 6. The Other Direction: From Standard Isoperimetry to Log-Sobolev
We now prove the other direction (i.e., the upper bound) in Theorem 1. Recall $\bar{\Lambda}$ defined around (1.5).
@@
Under Condition 1, for any $\alpha>0$,
(6.1) $\bar{\Lambda}(\alpha)$ $\leq\sqrt{\breve{\Theta}(\alpha)}.$
@@
We first prove the first inequality in (6.1). Let $A=A_{n}$ be a set approximately attaining $\Gamma_{n}(e^{-n\alpha},n\tau)$ (given in (3.1)) within a factor $1-\epsilon$ where $\epsilon>0$, i.e., $A$ has probability $e^{-n\alpha}$ and $\nu_{n}(A^{\sqrt{n\tau}})\geq(1-\epsilon)\Gamma_{n}(e^{-n\alpha},n\tau)$. Let $r:=\sqrt{n\tau}$. Since $f(t):=\ln\nu_{n}(A^{t})$ is monotone, it is differentiable almost everywhere on $(0,r)$. As a basic property of monotone functions, $f^{\prime}$ is integrable over $[0,r]$ and $\int_{0}^{r}f^{\prime}(t){\text{\rm d}}t\leq f(r)-f(0)$ (see e.g., [66, Corollary 4 on p.113]), which implies that there is some $r_{0}\in(0,r)$ such that $\frac{f(r)-f(0)}{r}\geq f^{\prime}(r_{0}),$i.e.,
$\frac{1}{r}\ln[\nu_{n}(A^{r})/\nu_{n}(A)]\geq\frac{\nu_{n}^{+}(A^{r_{0}})}{\nu_{n}(A^{r_{0}})}.$
@@
$\frac{1}{r}\ln[\nu_{n}(A^{r})/\nu_{n}(A)]\geq\frac{I_{n}(\nu_{n}(A^{r_{0}}))}{\nu_{n}(A^{r_{0}})}\geq\frac{I_{n}(\nu_{n}(A^{r}))}{\nu_{n}(A^{r})}.$
@@
(6.2) $\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$ $\geq\lim_{n\to\infty}\frac{\ln[\nu_{n}(A_{n}^{\sqrt{n\tau}})/\nu_{n}(A_{n})]}{n\sqrt{\tau}}\geq\limsup_{n\to\infty}\frac{I_{n}(a_{n,\tau})}{a_{n,\tau}\sqrt{n}},$
where $a_{n,\tau}=\nu_{n}(A_{n}^{\sqrt{n\tau}})$ and $\psi$ is defined in (3.12).
Since $A$ approximately attains $\Gamma_{n}(e^{-n\alpha},n\tau)$ within a factor $1-\epsilon$, by Statement 2 in Theorem 7 again, it holds that $\liminf_{n\to\infty}-\frac{1}{n}\ln a_{n,\tau}\geq\psi(\alpha,\tau).$ Hence, (6.2) implies that for any $\delta>0$, $\bar{\Lambda}(\psi(\alpha,\tau)-\delta)\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$ Taking $\tau\downarrow 0$ yields
(6.3) $\limsup_{\tau\downarrow 0}\bar{\Lambda}(\psi(\alpha,\tau)-\delta)\leq\xi(\alpha)$ $:=\limsup_{\tau\downarrow 0}\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$
We can assume that $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$, since otherwise, $\xi(\alpha)=\infty$ and nothing is needed to prove. Hence, given any $\delta>0$, we assume that $\psi(\alpha,\tau)>\alpha-\delta$ for all sufficiently small $\tau$, which, combined with the inequality above, implies
(6.4) $\bar{\Lambda}(\alpha-2\delta)\leq\xi(\alpha).$
We next prove $\xi(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$. Substituting the expression of $\psi(\alpha,\tau)$ into the bound above yields that
$\xi(\alpha)$ $=\limsup_{\tau\downarrow 0}\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$
@@
Since $\pi_{Y|W=i}=\pi_{X|W=i}$ is always a feasible solution to the supremum above, we have $\beta_{i}\leq\alpha_{i},i\in\{0,1\}$.
@@
We now need the HWI inequality for the Riemannian manifold setting [73, Corollary 20.13] [58].
Theorem 15 (HWI Inequality [73, Corollary 20.13]).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that Condition 1 holds. Then, for any two probability measures $\mu_{0}=\rho_{0}\nu,\mu_{1}=\rho_{1}\nu$ in $\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\mu_{1}\|\nu)<\infty$ and $\rho_{0}$ is Lipschitz, it holds that
@@
This theorem implies $|\partial^{-}D|(\mu\|\nu)\leq\sqrt{I(\mu\|\nu)}.$ In fact, this is an equality even for Polish metric probability measure space satisfying $\mathrm{\mathrm{C}D}(K,\infty)$; see [6, Theorem 9.3]. We restrict $\pi_{X|W=i},i\in\{0,1\}$ to be in $\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})$. Therefore,
$\xi(\alpha)$ $\leq\inf_{\begin{subarray}{c}\pi_{XW}:\pi_{X|W=i}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),\forall i,\\ D(\pi_{X|W}\|\nu|\pi_{W})=\alpha\end{subarray}}\sqrt{I(\pi_{X|W}\|\nu|\pi_{W})}=\sqrt{\breve{\Theta}(\alpha)}.$
Combining this with (6.4) yields that $\bar{\Lambda}(\alpha-2\delta)\leq\xi(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$ That is, for $\alpha,\delta>0$, $\bar{\Lambda}(\alpha)\leq\xi(\alpha+\delta)\leq\sqrt{\breve{\Theta}(\alpha+\delta)}.$ Since $\breve{\Theta}(t)$ is continuous in $t>0$, letting $\delta\downarrow 0$ yields $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$
@@
By an approximation argument, we may assume that the probability measures $\mu$ in the nonlinear log-Sobolev inequality have smooth density $f:={\text{\rm d}}\mu/{\text{\rm d}}\nu$. Obviously, $\int f\,d\nu=1$. Let $f^{\otimes n}$ be the $n$-fold product of $f$ with itself, which satisfies ${\text{\rm d}}\mu_{n}/{\text{\rm d}}\nu_{n}=f^{\otimes n}$ with $\mu_{n}=\mu^{\otimes n}$. Then, the coarea formula (see e.g., [25, Exercise III.12]) states:
@@
$\Theta_{k}(\alpha):=\inf_{\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}):D(\mu\|\nu^{\otimes k})\geq\alpha}I(\mu\|\nu^{\otimes k}).$
@@
Appendix A: Proof of Statement 2 of Theorem 1
Our proof for $\Lambda(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$ still works. We now show that if Condition 2 does not hold, then $\breve{\Theta}(\alpha)=0$ for any $\alpha>0$. We prove this by contradiction.
Suppose that $\breve{\Theta}(\alpha_{0})>0$ for some $\alpha_{0}>0$. Then, $\breve{\Theta}(\alpha)\geq C\alpha$ for $\alpha\geq\alpha_{0}$ where $C=\frac{\breve{\Theta}(\alpha_{0})}{\alpha_{0}}>0$. Thus,
(6.14) $C(D(\mu\|\nu)-\alpha_{0})\leq I(\mu\|\nu),\;\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}).$
Equivalently, for any smooth function $f\in L^{2}(\mathcal{M},\nu)$:
$\int f^{2}\ln f^{2}{\text{\rm d}}\nu-\left(\int f^{2}{\text{\rm d}}\nu\right)\ln\left(\int f^{2}{\text{\rm d}}\nu\right)-\alpha_{0}\int f^{2}{\text{\rm d}}\nu\leq\frac{4}{C}\int|\nabla f|^{2}{\text{\rm d}}\nu$
We now use Herbst’s argument to derive exponential integrability from this log-Sobolev inequality. Let $F(x)=d(x,x_{0})$ be the distance function. By Rademacher’s theorem, $|\nabla F|\leq 1$ almost everywhere. Define the moment generating function for $F$: $\psi(\lambda)=\int e^{\lambda F}{\text{\rm d}}\nu$. We assume $\psi(\lambda)<\infty$ for all $\lambda\geq 0$.
Apply the LSI to the function $f^{2}=e^{\lambda F}$. Then $|\nabla f|^{2}=\frac{\lambda^{2}}{4}e^{\lambda F}|\nabla F|^{2}$. Substituting this into the LSI:
$\int\lambda Fe^{\lambda F}{\text{\rm d}}\nu-\psi(\lambda)\ln\psi(\lambda)-\alpha_{0}\psi(\lambda)\leq\frac{4}{C}\int\frac{\lambda^{2}}{4}e^{\lambda F}|\nabla F|^{2}{\text{\rm d}}\nu.$
Since $|\nabla F|\leq 1$, we have:
$\frac{\lambda\psi^{\prime}(\lambda)-\psi(\lambda)\ln\psi(\lambda)-\alpha_{0}\psi(\lambda)}{\lambda^{2}\psi(\lambda)}\leq\frac{1}{C}.$
Notice that the left-hand side is the derivative of the function $\lambda\mapsto\frac{\ln\psi(\lambda)+\alpha_{0}}{\lambda}$, i.e., $\frac{{\text{\rm d}}}{{\text{\rm d}}\lambda}\left(\frac{\ln\psi(\lambda)+\alpha_{0}}{\lambda}\right)\leq\frac{1}{C}$. Integrating from a constant $\lambda_{0}$ to $\lambda$ yields $\frac{\ln\psi(\lambda)+\alpha_{0}}{\lambda}-C_{1}\leq\frac{\lambda}{C}$ where $C_{1}=\frac{\ln\psi(\lambda_{0})+\alpha_{0}}{\lambda_{0}}$. Thus,
$\psi(\lambda)\leq e^{\lambda^{2}/C+C_{1}\lambda-\alpha_{0}}.$
Note that in the proof above, we assume $\psi(\lambda)<\infty$ for all $\lambda\geq 0$. This is in fact unnecessary since we can obtain the same inequality by substituting $F\leftarrow F_{n}=\min\{F,n\}$ in the proof above and letting $n\to\infty$.
Using Markov’s inequality, we obtain the tail estimate for any $r>0$:
$\nu(F\geq r)=\nu(e^{\lambda F}\geq e^{\lambda r})\leq\frac{\mathbb{E}[e^{\lambda F}]}{e^{\lambda r}}\leq e^{\lambda^{2}/C+C_{1}\lambda-\alpha_{0}-\lambda r}.$
To get the tightest bound, we minimize the exponent with respect to $\lambda$. Setting $\lambda=C(r-C_{1})/2$ gives $\nu(F\geq r)\leq e^{-\frac{C(r-C_{1})^{2}}{4}-\alpha_{0}}$.
We now use this tail estimate to evaluate $\mathbb{E}_{\nu}[e^{\alpha F^{2}}]$. We use the layer-cake representation $\mathbb{E}_{\nu}[e^{\alpha F^{2}}]=\int_{0}^{\infty}\nu(e^{\alpha F^{2}}>t){\text{\rm d}}t$. Let $t=e^{\alpha r^{2}}$, which implies $r=\sqrt{\frac{\ln t}{\alpha}}$. Changing variables $\mathbb{E}_{\nu}[e^{\alpha F^{2}}]=1+\int_{0}^{\infty}\nu(F>r)\cdot 2\alpha re^{\alpha r^{2}}{\text{\rm d}}r$. Substitute the tail bound into the integral:
$\mathbb{E}_{\nu}[e^{\alpha F^{2}}]\leq 1+\int_{0}^{\infty}2e^{-\frac{C(r-C_{1})^{2}}{4}-\alpha_{0}}\cdot 2\alpha re^{\alpha r^{2}}{\text{\rm d}}r=1+4\alpha\int_{0}^{\infty}re^{\alpha r^{2}-\frac{C(r-C_{1})^{2}}{4}-\alpha_{0}}{\text{\rm d}}r.$
Obviously, the integral converges to a finite value for any $\alpha<\frac{C}{4}$. This contradicts with that Condition 2 does not hold. Hence, $\breve{\Theta}(\alpha)=0$ for any $\alpha>0$.
Appendix B: Proof of Theorem 2
Observe that $\{\nu_{m}^{\otimes n}\}$ converges to $\nu^{\otimes n}$ in total-variation distance and in addition $\nu_{m}^{\otimes n}(B)\geq(1-1/m)^{n}\nu^{\otimes n}(B)$ for any Borel set $B$. E. Milman [51, Theorem 6.10] showed that in this case, for all $a\in[0,1]$, $I_{\nu^{\otimes n}}(a)=\lim_{m\to\infty}I_{\nu_{m}^{\otimes n}}(a),$ and consequently, $I_{\nu^{\otimes n}}$ is concave. Let $A$ be an isoperimetric minimizer for $\nu_{m}^{\otimes n}$ in the standard sense. By Theorem 12, for any $r>0$,
$\frac{I_{\nu_{m}^{\otimes n}}(a)}{a}\geq\frac{1}{r}\ln\frac{\nu_{m}^{\otimes n}(A^{r})}{a}\geq\frac{1}{r}\ln\frac{(1-1/m)^{n}\nu^{\otimes n}(A^{r})}{a}\geq\frac{1}{r}\ln\frac{(1-1/m)^{n}\Gamma_{\nu^{\otimes n}}(a,r)}{a}.$
Taking the limit as $m\to\infty$, we obtain $\frac{I_{\nu^{\otimes n}}(a)}{a}\geq\frac{1}{r}\ln\frac{\Gamma_{\nu^{\otimes n}}(a,r)}{a}.$ Note that $\mathrm{RCD}(0,\infty)$ is satisfied by $\nu$, and thus, if Condition 2 is additionally satisfied, then Theorem 8 is valid for $\nu$. The remaining part of the proof is exactly the same as Theorem 1.
As for the other direction, noting that $I_{\nu^{\otimes n}}$ is concave for every $n$ and $(\mathcal{M},d,\nu)$ is a $\mathrm{RCD}(0,\infty)$-space, our first proof given in Section 6 still works for this space, but with the HWI inequality replaced by [6, Theorem 9.3]. Hence, $\Lambda=\sqrt{\breve{\Theta}}$ holds.
If Condition 2 is not satisfied, then $\Lambda\leq\sqrt{\breve{\Theta}}$ still holds, and thus, by the same proof in Appendix A, $\Lambda=\sqrt{\breve{\Theta}}=0$.
Appendix C: Proof of Theorem 3
@@
Appendix C-1. Inequality (b)
We first prove Inequality (b) in Theorem 3. Due to Theorem 11, or more precisely, due to (3.1) and (4.2) in [52] but with $K_{CD}\geq 0$, it holds that for any $n$,
@@ Appendix C-1. Inequality (b)
Appendix C-2. Inequality (k)
Inequality (k) is exactly the following inequality: for any $n\geq 1$,